# Joint Channel, Data and Radar Parameter Estimation for AFDM Systems in Doubly-Dispersive Channels

### v2

@@ II-B OFDM Signal Model
$\mathbf{s}_{\text{OFDM}}=\mathbf{F}_{N}^{\mathsf{H}}\cdot\mathbf{x}\in\mathbb{C}^{N\times 1},$ (8)
@@
$\bm{\varPhi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\pi(N^{2}-2N\ell_{p})},e^{-j2\pi(N^{2}-2N(\ell_{p}-1))},\dots,e^{-j2\pi(N^{2}-2N)}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-1ex}$ (22)
@@
$\mathbf{G}_{\text{OFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot\mathbf{F}_{N}\cdot(\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot\mathbf{F}_{N}^{\mathsf{H}}.$ (12)
@@ II-C OTFS Signal Model
Consider an OTFS system in which an input information matrix $\mathbf{X}\in\mathcal{C}^{K\times M}$, containing $K\times M$ complex symbols taken from an arbitrary constellation $\mathcal{C}$, is modulated into the transmit signal [^7] For simplicity, we assume that all pulse-shaping operations are perfect ($i.e.,$ the corresponding pulse-shaping waveforms are rectangular), such that the corresponding sample matrices can be reduced to identity matrices.
@@
We highlight that the notation in equation (13) is in line with the strategy described $e.g.$ in [34], whereby the OTFS signals are first vectorized and then appended with a CP of length $N_{CP}$ in order to eliminate inter-frame interference, in similarity with OFDM. Taking advantage of such similarity, and in order to allow for direct comparisons between the two waveforms, we shall hereafter set $K\times M=N$.
@@
$\mathbf{G}_{\text{OTFS}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{F}_{M}\otimes\mathbf{I}_{K})\cdot(\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{F}_{M}^{\mathsf{H}}\otimes\mathbf{I}_{K}).$ (19)
@@ II-D AFDM Signal Model
It was shown in [13] that after going through a doubly-dispersive channel, an AFDM modulated symbol vector as given in equation (20) with the inclusion of a CPP as described in [19], can be modeled similarly to equation (4), only with the CP matrix $\bm{\Phi}_{p}$ of equation (5) replaced by the corresponding CPP matrix $\bm{\varPhi}_{p}$ as given in equation (22), such that the corresponding received signal is given by
@@
$\mathbf{G}_{\text{AFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{\Lambda}_{2}\mathbf{F}_{N}\mathbf{\Lambda}_{1})\cdot(\bm{\varPhi}_{p}^{c_{1}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{\Lambda}_{1}^{\mathsf{H}}\mathbf{F}_{N}^{\mathsf{H}}\mathbf{\Lambda}_{2}^{\mathsf{H}}),$ (26)
@@ III Joint Channel and Data Estimation
where the matrices $\mathbf{\Gamma}_{p}$ captures the long-term delay-Doppler statistics of the channel, which are assumed to be known.
From equation (27), it is evident that each received signal in $\mathbf{y}$ can described by
@@ III-A JCDE via Parametric Bilinear Gaussian Belief Propagation
For starters, referring to Figure 2, it is initially assumed [^10] Later we shall adopt the strategy proposed in [19] and consider, for the AFDM case, the alternative of utilizing only one pilot followed by $B-1$ zeros among the $N$ symbols in $\mathbf{x}$, for the purpose of enabling the RPE scheme of Section IV-B. that in any of the three systems compared, a block of $B$ pilots [^11] We use the first column of the $B\times B$ Zadoff-chu sequence [1] as our pilot column. is embedded in the $N\times 1$ symbol vector $\mathbf{x}$.
@@ III-A1 Soft IC
where $\mathbb{E}_{x}$ refers to expectation over the all possible symbols $x$ in the constellation $\mathcal{C}$, while $\mathbb{E}_{h}$ refers to expectation over all possible outcomes of $h\sim\mathcal{CN}(0,\sigma^{2}_{h})$, respectively [^12] Although the average power of each channel path may be generally different, especially in sub-6GHz channels where the propagation lengths of different paths can vary from meters to kilometers, it is commonly assumed in high-frequency (mmWave and TeraHz) channels, where ranges are much shorter, that the gains of all paths follow the same distribution [18, 19, 20]..
@@
while $\tilde{\sigma}^{2(i)}_{x:{m,n}}$ are the soft IC conditional variances given by
$\tilde{\sigma}^{2(i)}_{x:{m,n}}\triangleq\mathbb{E}_{x\neq x_{m},h}\big{[}|\tilde{y}_{x:{m,n}}^{(i)}-\tilde{\gamma}_{x:{m,n}}^{(i)}x_{m}|^{2}\big{]}$
@@ III-A2 Belief Generation
Then, the belief corresponding to the $m$-th symbol $x_{m}$ at the $n$-th factor node is obtained by combining the contributions of all signals in $\mathbf{y}$, excluding $y_{n}$, under the PDF
$\!\!\!\!p_{\text{x}|x_{m}}(\text{x}|x_{m})\!=\!\prod_{q\neq n}^{N}\!p_{\text{y}_{q}|\text{x}_{m}}(\text{y}_{q}|x_{m})\propto\text{exp}\bigg{(}\!\!\!-\!\frac{|\text{x}\!-\!\tilde{x}_{n,m}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}}\!\bigg{)},\!$ (38a)
@@
Similarly, the extrinsic beliefs of the channel gains can be shown to follow the approximate distribution
@@ III-A3 Soft RG
where $p_{\text{x}|x}\big{(}\text{x}|x;\tilde{x}_{n,m}^{(i)},\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\big{)}$ and $p_{\text{h}|h}\big{(}\text{h}|h;\tilde{h}_{n,p}^{(i)},\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}\big{)}$ are the likelihood functions of the data and channel beliefs given in equations (38a) and (40), respectively, only with the notation extended to explicitly show the parameters (means and variances) computed above.
After obtaining $\hat{x}_{n,m}^{(i)}$ as per equation (44), the final output is computed by damping the result in order to improve convergence [38], yielding
@@
which are respectively damped via
$\hat{h}_{n,p}^{(i)}=\beta_{h}\hat{h}_{n,p}^{(i)}+(1-\beta_{h})\hat{h}_{n,p}^{(i-1)}$ (49)
@@
$\sigma_{h:{n,p}}^{2(i)}=\beta_{h}\sigma_{h:{n,p}}^{2(i)}+(1-\beta_{h})\sigma_{h:{n,p}}^{2(i-1)}.$ (50)
@@
III-B Performance Analysis
@@
First, we compare in Figure 4 the communications and channel estimation performances, in terms of BER and normalized mean square error (NMSE), respectively, achieved by PBiGaBP employed over OTFS, OFDM and AFDM waveforms. In this comparison, all systems employ an equal number $N_{p}$ of pilot symbols, although each system has its own structure (see Section II and Figure 2). The results show that PBiGaBP-based JCDE with AFDM and OTFS outperforms a system employing the same proposed technique but employing a legacy OFDM waveform, both in terms of BER and NMSE, with a slight advantage of AFDM over the OTFS alternative.
We highlight that a comparison between the proposed JCDE algorithm and other techniques from related literature is not offered in Figure 4, as we are unaware of any alternative JCDE solution for AFDM systems. Therefore, Figure 4 compares the proposed PBiGaBP-based JCDE against the GaBP alternative to the ML-based AFDM channel estimation method from [20].
We emphasize, however, that the technique in [20] is not a JCDE scheme, being rather strictly pilot-based, such that the latter SotA method enjoys, in this comparison, the advantage of not dealing with the additional burden of performing data detection, which the proposed algorithm does. In addition, as indicated in the legend of the figure, the comparison awards the SotA approach a power advantage, by means of allowing the latter to employ pilot signals with higher power than that used in the proposed method.
The results show that the proposed PBiGaBP-based JCDE scheme outperforms the GaBP-based method in [20] even when the latter employs a pilot with 10 times as higher power, which is a consequence of the fact that the proposed scheme takes advantage not only of the pilot, but also payload symbols to estimate the channel. The results also motivates us hereafter to no longer consider OFDM, as it is clear that this waveform is outperformed by OTFS and AFDM in doubly-dispersive channels.
@@
Figure 4: BER Performance of single pilot AFDM at the UE via the proposed PBiGaBP algorithm compared to the SotA methods.
@@ IV Radar Parameter Estimation in Doubly Dispersive Channels
$\mathbf{h}\!\triangleq\![h_{0,0},\dots,h_{0,D_{\nu}\!-\!1},\dots,h_{K_{\tau}\!-\!1,0},\dots,h_{K_{\tau}\!-\!1,D_{\nu}\!-\!1}]^{\mathsf{T}}\!\!\!.\vspace{-1ex}$ (54c)
@@
Monostatic RPE: A first scenario is the classic monostatic case, in which the BS processes round-trip signals reflected by surrounding scatters (targets) in order to extract their distances and velocities. Here, a main practical challenge is the self-interference caused by signals transmitted by the BS at the same time when reflected signals return. The issue is often ignored in related literature on OFDM- and OTFS-based monostatic RPE [8, 18], implying a strong and not entirely practical assumption that self-interference is perfectly avoided by FD technologies. To emphasize this point, it has been shown that AFDM has an advantage over OFDM and OTFS as it allows for the concentration of a single pilot symbol within a transmission block [^15] Observe from equations (25) and (26) that unlike OFDM and OTFS, the AFDM receive signal occupies the entire time-frequency span even when a single entry in the transmit vector $\mathbf{x}$ is non-zero, such that concentrating piloting to a single symbol does not sacrifice the DoF of the estimator., as illustrated in Figure 2, with the corresponding interference mitigated via a de-chirping operation [20]. We will therefore focus on this case.
@@
User-centric RPE: In this case, the BS first transmits a pilot symbol vector used by the UE to perform RPE, and subsequently transmits data from which the UE proceeds to perform JCDE as described in Section III in order to track the channel and decode the information. Under the assumption that delays and Dopplers change at a rate slower than $h_{p}$ [45], the alternation of JCDE and RPE stages would enable the system to minimize the overhead associated with frequent pilot transmissions. Due to the lack of space, this scenario will not be further developed here, but we have shown in [27] that a well-designed JCDE algorithm is capable of tracking time-varying channels even without periodic pilot transmissions, as long as the BER of the system is maintained sufficiently low, $e.g.$ via adequate channel coding. It is not difficult to envision that by combining the contribution of this article with the one in [27], the user-centric RPE scenario can be addressed.
@@
Bistatic RPE: This refers to the case when multiple (passive) BSs perform RPE using signals transmitted by one (active) BS. The scenario can be similar either to the latter user-centric case, if the passive BSs have no knowledge of the transmit vector $\mathbf{x}$, or to the monostatic case, if $\mathbf{x}$ can be assumed known to all passive BSs. It is clear, therefore, that this scenario is a generalization of both, which is best addressed in a dedicated follow up article, especially since the latter assumption requires the BSs to be interconnected via a fronthaul network, which in turn might be subjected to impairments such as delays, noise, capacity limitations etc.

### v4

@@ II-B OFDM Signal Model
$\mathbf{s}_{\text{OFDM}}=\mathbf{F}_{N}^{\mathsf{H}}\cdot\mathbf{x}\in\mathbb{C}^{N\times 1},\vspace{-0.5ex}$ (8)
@@
$\mathbf{G}_{\text{OFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot\mathbf{F}_{N}\cdot(\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot\mathbf{F}_{N}^{\mathsf{H}},$ (12)
with the CP phase matrices $\mathbf{\Phi}_{p}$ appearing in equation (4) reduced to identity matrices [16]; $i.e.,$ $\phi_{\mathrm{CP}}(n)=0$ in equation (5) since there is no phase offset in the CP of OFDM signals.
@@ II-C OTFS Signal Model
Consider an OTFS system in which an input information matrix $\mathbf{X}\in\mathcal{C}^{K\times M}$, containing $K\times M$ complex symbols taken from an arbitrary constellation $\mathcal{C}$, is modulated into the transmit signal [^8] For simplicity, we assume that all pulse-shaping operations utilize rectangular waveforms, such that the corresponding sample matrices can be reduced to identity matrices.
@@
We highlight that the notation in equation (13) is in line with the strategy described $e.g.$ in [40], whereby the OTFS signals are first vectorized and then appended with a CP of length $N_{\mathrm{CP}}$ in order to eliminate inter-frame interference, in similarity with OFDM. Taking advantage of such similarity, and in order to allow for direct comparisons between the two waveforms, we shall hereafter set $K\times M=N$.
@@
$\mathbf{G}_{\text{OTFS}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{F}_{M}\otimes\mathbf{I}_{K})\cdot(\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{F}_{M}^{\mathsf{H}}\otimes\mathbf{I}_{K}),$ (19)
in which, similar to the OFDM waveform, the CP phase matrices $\mathbf{\Phi}_{p}$ have been reduced to identity matrices [16].
@@ II-D AFDM Signal Model
$\bm{\varPhi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\pi c_{1}(N^{2}-2N\ell_{p})},e^{-j2\pi c_{1}(N^{2}-2N(\ell_{p}-1))},\dots,e^{-j2\pi c_{1}(N^{2}-2N)}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-1ex}$ (22)
@@
It was shown in [16] that after going through a doubly-dispersive channel, an AFDM modulated symbol vector as given in equation (20) with the inclusion of a CPP as described in [22], can be modeled similarly to equation (4), only with the CP matrix $\bm{\Phi}_{p}$ of equation (5) replaced by the corresponding CPP matrix $\bm{\varPhi}_{p}$ as explicitly given in equation (22) ($i.e.,$ $\phi_{\mathrm{CP}}(n)=c_{1}(N^{2}-2Nn)$ in equation (5)), such that the corresponding received signal is given by
@@
$\mathbf{G}_{\text{AFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{\Lambda}_{2}\mathbf{F}_{N}\mathbf{\Lambda}_{1})\cdot(\bm{\varPhi}_{p}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{\Lambda}_{1}^{\mathsf{H}}\mathbf{F}_{N}^{\mathsf{H}}\mathbf{\Lambda}_{2}^{\mathsf{H}}),$ (26)
@@ III Joint Channel and Data Estimation
where the matrices $\mathbf{\Gamma}_{p}$ captures the long-term delay-Doppler statistics of the channel, which are assumed to be known [^11] According to [41, 42, 43], the delay and Doppler shifts can be assumed constant during multiple frame transmissions, enabling sporadic estimation of $\mathbf{\Gamma}_{p},\forall p$ via RPE schemes utilizing pilots such as the method described in the latter part of Section IV-B. However, the addition of this procedure into the JCDE scheme transforms this into an RPE-aided JCDE method ($i.e.,$ user-centric RPE) which while consolidating and combining the seperate aspects of ISAC, deserves further attention and will be addressed in a follow-up article..
From equation (27), it is evident that each received signal in $\mathbf{y}$ can be described by
@@ III-A JCDE via Parametric Bilinear Gaussian Belief Propagation
For starters, referring to Figure 2, it is initially assumed [^12] Later we shall adopt the strategy proposed in [22] and consider, for the AFDM case, the alternative of utilizing only one pilot followed by a null-guard interval of length $B-1$ among the $N$ symbols in $\mathbf{x}$, for the purpose of enabling the RPE scheme of Section IV-B. that in any of the three systems compared, a block of $B$ pilots [^13] We use the first column of the $B\times B$ Zadoff-Chu sequence [1] as our pilot column. is embedded in the $N\times 1$ symbol vector $\mathbf{x}$.
@@ III-A1 Soft IC
where $\mathbb{E}_{x}$ refers to expectation over the all possible symbols $x$ in the constellation $\mathcal{C}$, while $\mathbb{E}_{h}$ refers to expectation over all possible outcomes of $h\sim\mathcal{CN}(0,\sigma^{2}_{h})$, respectively [^14] Although the average power of each channel path may be generally different, especially in sub-6GHz channels where the propagation lengths of different paths can vary from meters to kilometers, it is commonly assumed in high-frequency (mmWave and Terahertz) channels, where ranges are much shorter, that the gains of all paths follow the same distribution [21, 22, 23]..
@@
while $\tilde{\sigma}^{2(i)}_{x:{m,n}}$ are the soft IC conditional variances approximated by replacing the instantaneous values with the long-term statistics [47, 48] as
$\tilde{\sigma}^{2(i)}_{x:{m,n}}\triangleq\mathbb{E}_{x,h}\big{[}|\tilde{y}_{x:{m,n}}^{(i)}-\tilde{\gamma}_{x:{m,n}}^{(i)}x_{m}|^{2}\big{]}$
@@ III-A2 Belief Generation
Then, as a result of SGA, the belief corresponding to the $m$-th symbol $x_{m}$ at the $n$-th factor node is obtained by combining the contributions of all signals in $\mathbf{y}$, excluding $y_{n}$, under the PDF
$\!\!\!\!p_{\text{x}|x_{m}}(\text{x}|x_{m})\!=\!\prod_{q\neq n}^{N}\!p_{\text{y}_{q}|\text{x}_{m}}(\text{y}_{q}|x_{m})\propto\text{exp}\bigg{(}\!\!\!-\!\frac{|\text{x}\!-\!\tilde{x}_{n,m}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}}\!\bigg{)},\!$ (38)
@@
Similarly, under SGA, the extrinsic beliefs of the channel gains can be shown to follow the approximate distribution
@@ III-A3 Soft RG
where $p_{\text{x}|x}\big{(}\text{x}|x;\tilde{x}_{n,m}^{(i)},\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\big{)}$ and $p_{\text{h}|h}\big{(}\text{h}|h;\tilde{h}_{n,p}^{(i)},\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}\big{)}$ are the likelihood functions of the data and channel beliefs given in equations (38) and (40), respectively, only with the notation extended to explicitly show the parameters (means and variances) computed above.
@@
After obtaining $\hat{x}_{n,m}^{(i)}$ as per equation (44), the final output is computed by damping the result with a damping factor $10<\beta_{x}<$ 1 in order to improve convergence [49], yielding
@@
which are then damped with a damping factor $10<\beta_{h}<$ 1 as
$\hat{h}_{n,p}^{(i)}=\beta_{h}\hat{h}_{n,p}^{(i)}+(1-\beta_{h})\hat{h}_{n,p}^{(i-1)},$ (49)
@@
$\hat{\sigma}_{h:{n,p}}^{2(i)}=\beta_{h}\hat{\sigma}_{h:{n,p}}^{2(i)}+(1-\beta_{h})\hat{\sigma}_{h:{n,p}}^{2(i-1)}.$ (50)
@@
Remark 1: Although the proposed PBiGaBP-based JCDE algorithm is designed based on the Bayes-optimal GaBP framework, to the best of our knowledge, there exists no asymptotic analysis of the large-system limit for iterative convergence analysis of Bayesian parametric bilinear inference algorithms. While semi-analytical methods have been proposed for the prediction of the number of iterations required for fixed-point convergence in large scale systems [29, 52], they cannot be used under correlated observations. Therefore, for the performance analysis detailed in Section III-C, rigorous numerical simulations are provided to demonstrate the superiority of the proposed technique.
III-B Complexity Analysis
Since there are no high-complexity matrix inversions in the proposed PBiGaBP-based JCDE scheme detailed in Algorithm 1, the order of complexity strictly depends on the scalar operations executed. Therefore, the computational complexity of the proposed method is $\mathcal{O}(N^{2}(P+1))$ per iteration, which is linear in terms of the operations performed, and of a much lower complexity than SotA methods based on linear minimum mean square error (LMMSE) filtering, which incur a cost of $\mathcal{O}(N^{3})$ due to a computation of LMMSE filter involving a matrix inversion operation.
The proposed method also has comparable complexity to the SotA parametric bilinear generalized approximate message passing (PBiGAMP) scheme given in [1], which also is of order $\mathcal{O}(N^{2}(P+1))$, but far outperforms the latter method in terms of performance, therefore offering an excellent performance-complexity tradeoff.
III-C Performance Analysis
@@
First, using the Linear GaBP  [^18] Linear GaBP is a technique that uses GaBP for both channel estimation (under full knowledge of the frame) and data decoding (under full knowledge of the channel coefficients), combined with the use of an LMMSE filtering-based initialization for channel estimation. The approach provides a lower bound on the performance for the proposed method in exchange for a large computational cost. as a baseline, we compare in Figure 3 the communications and channel estimation performances, in terms of BER and normalized mean square error (NMSE), respectively, achieved by PBiGaBP employed over OTFS, OFDM and AFDM waveforms. In this comparison, all systems employ an equal number $N_{p}$ of pilot symbols, although each system has its own structure (see Section II and Figure 2).
@@
The results show that PBiGaBP-based JCDE with AFDM and OTFS outperforms a system employing the same proposed technique but employing a legacy OFDM waveform, both in terms of BER and NMSE, with a slight advantage of AFDM over the OTFS alternative.
We highlight that a comparison between the proposed JCDE algorithm and other techniques from related literature is not offered in Figure 3, as we are unaware of any alternative JCDE solution for AFDM systems. Therefore, Figure 4 compares the proposed PBiGaBP-based JCDE both in terms of BER and NMSE performance against the GaBP alternative to the ML-based AFDM channel estimation method from [23], where channel estimation is done solely based on the single pilot via GaBP with an accompanying GaBP-based data decoding.
We emphasize, however, that the technique in [23] is not a JCDE scheme, being rather strictly pilot-based like traditional channel estimation, such that in comparison, the latter SotA method enjoys the advantage of not dealing with the additional burden of performing data detection, which the proposed algorithm does. In addition, as indicated in the legend of the figure, the comparison awards the SotA approach a power advantage by allowing the latter to employ pilot signals with higher power than those used in the proposed method.
(a) BER Performance.
(b) NMSE Performance.
Figure 4: Performance of the single pilot QPSK-AFDM system at the UE via the proposed PBiGaBP algorithm compared to the SotA methods.
The results show that the proposed PBiGaBP-based JCDE scheme outperforms the traditional pilot-only GaBP-based method for the single pilot AFDM frame structure proposed in [23] and illustrated in Figure 2, even when the latter employs a pilot with 10 times as higher power, which is a consequence of the fact that the proposed scheme takes advantage not only of the pilot, but also payload symbols ($i.e.,$ soft pilots) to estimate the channel. The results also motivates us hereafter to no longer consider OFDM, as it is clear that it is outperformed by OTFS and AFDM in doubly-dispersive channels.
@@ IV Radar Parameter Estimation in Doubly Dispersive Channels
$\mathbf{h}\!\triangleq\![h_{0,0},\dots,h_{0,D_{\nu}\!-\!1},\dots,h_{K_{\tau}\!-\!1,0},\dots,h_{K_{\tau}\!-\!1,D_{\nu}\!-\!1}]^{\mathsf{T}}\!\!\!.$ (54c)
@@
Monostatic RPE: A first scenario is the classic monostatic case, in which the BS processes round-trip signals reflected by surrounding scatters (targets) in order to extract their distances and velocities. Here, a main practical challenge is the self-interference caused by signals transmitted by the BS at the same time when reflected signals return. The issue is often ignored in related literature on OFDM- and OTFS-based monostatic RPE [11, 21], implying a strong and not entirely practical assumption that self-interference is perfectly avoided by FD technologies. To emphasize this point, it has been shown that AFDM has an advantage over OFDM and OTFS as it allows for the concentration of a single pilot symbol within a transmission block [^20] Observe from equations (25) and (26) that unlike OFDM and OTFS, the AFDM receive signal occupies the entire time-frequency span even when a single entry in the transmit vector $\mathbf{x}$ is non-zero, such that concentrating piloting to a single symbol does not sacrifice the DoF of the estimator., as illustrated in Figure 2, with the corresponding self-interference mitigated via a simple analog dechirping and filtering operation [23]. We will therefore focus on this case.
@@
User-centric RPE: In this case, the BS first transmits a pilot symbol vector used by the UE to perform RPE, and subsequently transmits data from which the UE proceeds to perform JCDE as described in Section III in order to track the channel and decode the information. Under the assumption that delay and Doppler shifts change at a rate slower than $h_{p}$ [57], the alternation of JCDE and RPE stages would enable the system to minimize the overhead associated with frequent pilot transmissions. Due to the lack of space, this scenario will not be further developed here, but we have shown in [28] that a well-designed JCDE algorithm is capable of tracking time-varying channels even without periodic pilot transmissions, as long as the BER of the system is maintained sufficiently low, $e.g.$ via adequate channel coding. It is not difficult to envision that by combining the contribution of this article with the one in [28], the user-centric RPE scenario can be addressed.
@@
Bistatic RPE: This refers to the case when passive receivers – could be BSs or UEs– which are placed at different locations with respect to the actively transmitting BS perform RPE using the received signals from the active BS [58, 59, 60, 61]. The scenario can be similar either to the latter user-centric case, if the passive receiver have no knowledge of the transmit vector $\mathbf{x}$ [61], or to the monostatic case, if $\mathbf{x}$ can be assumed known via pilot symbols, a preamble or a reference signal [58, 59, 60] to all the passive receivers. It is clear, therefore, that this scenario is a generalization of both, which is best addressed in a dedicated follow-up article, especially since the latter assumption – if BSs are considered – requires them to be interconnected via a fronthaul network, which in turn might be subjected to impairments such as delays, noise, capacity limitations etc.
For the rest of this manuscript, wlg and due to space limitations, we will focus on monostatic RPE.